OpenStudy (anonymous):

(Sqrt(x-4)+1) = sqrt(x+9)

OpenStudy (anonymous):

\[\LARGE \sqrt{x-4}+1=\sqrt{x+9}\] squaring both sides we have: \[\LARGE (\sqrt{x-4}+1)^2=(\sqrt{x+9})^2\] now use this formula: \[\LARGE (a+b)^2=a^2+2ab+b^2 \] for the left side... and for the right side just cancel out square root with square

OpenStudy (anonymous):

so would it be x^2-8+1=x+9

OpenStudy (anonymous):

no , try again.

OpenStudy (anonymous):

or show what have you done , so we can correct you.

OpenStudy (anonymous):

i tried to use the formula for the left side

OpenStudy (anonymous):

can I see how did you use it?

OpenStudy (anonymous):

i was just trying to see if thats how it would come out when u plug the numbers in for a b and c in the formulat

OpenStudy (anonymous):

there's no C ... I'll ask you again. Can I see how did you use it?

OpenStudy (anonymous):

\[x^2+2(1)(-4)+1^2\]

OpenStudy (anonymous):

\[\Large (\sqrt{x-4}+1)^2=(\sqrt{x-4})^2+2\cdot[ 1\cdot \sqrt{x-4}]+1^2 \]

OpenStudy (anonymous):

ok i see what i did wrong but now what do you do with all of this?

OpenStudy (anonymous):

to simplify ... \[\LARGE (\sqrt{x-4})^2+2\sqrt{x-4}+1=\] \[\LARGE x-4+1+2\sqrt{x-4 }=\] \[\LARGE x-3+2\sqrt{x-4}\] so your equation would look like: \[\LARGE x-3+2\sqrt{x-4}=x+9 \] ... what can we do now? :)

OpenStudy (anonymous):

cancel the x's and subtract the -3?

OpenStudy (anonymous):

yes... \[\LARGE 2\sqrt{x-4}=9+3 \] \[\LARGE 2\sqrt{x-4}=12 \] what can we do now?

OpenStudy (anonymous):

im not exactly sure..... do you have to square it?

OpenStudy (anonymous):

not actually :) (we can do it now, but it just would give us bigger numbers) we want to simplify it a bit :) let's divide both sides by 2 and we get?

OpenStudy (anonymous):

sqrt(x-4) =6

OpenStudy (anonymous):

well done... now we square both sides. and we get?

OpenStudy (anonymous):

x-4=36

OpenStudy (anonymous):

x=40

OpenStudy (anonymous):

That's correct. Well done . ;)

OpenStudy (anonymous):

thank you